Kiley Corporation had the following data for the most recent year (in millions). The new CFO believes (1) that an improved inventory management system could lower the average inventory by $4,000, (2) that improvements in the credit department could reduce receivables by $2,000, and (3) that the purchasing department could negotiate better credit terms and thereby increase accounts payable by $2,000. Furthermore, she thinks that these changes would not affect either sales or the costs of goods sold. If these changes were made, by how many days would the cash conversion cycle be lowered?

 

Original

Revised

Annual sales: unchanged

$110,000

$110,000

Cost of goods sold: unchanged

$80,000

$80,000

Average inventory: lowered by $4,000

$20,000

$16,000

Average receivables: lowered by $2,000

$16,000

$14,000

Average payables: increased by $2,000

$10,000

$12,000

Days in year

365

365

 

a. 49.8

b. 41.2

c. 34.0

d. 37.4

e. 45.3
Answer #1

Solution-

Days Inventory outstanding= average inventory / Cost of goods sold*365

Original =20000/80000*365= 91.25 days

Revised =16000/80000*365= 73 days

Days sales outstanding = average receivables/ annual sales*365

Original= 16000/110000*365=53.09 days

Revised = 14000/110000*365= 46.45 days

Days payable outstanding = average payable/cost of goods sold*365

Original = 10000/80000*365=45.62 days

Revised =12000/80000*365= 54.75 days

Cash conversion cycle= days inventory outstanding+ days sales outstanding- days payable outstanding

Original= 91.25 days+53.09 date- 45.62 days=98.72 days

Revised= 73 days +46.45 days -54.75 days=64.70 days

Net effect = Original - Revised

=98.72-64.70=34.02 days

So correct answer is option C i.e. 34.0 days